# Towards scalable quantum computations of atomic nuclei

### v1

@@ Abstract
We solve the nuclear two-body and three-body bound states via quantum simulations of pionless effective field theory on a lattice in position space. While the employed lattice remains small, the usage of local Hamiltonians including two- and three-body forces ensures that the number of Pauli terms scales linearly with increasing numbers of lattice sites. We use an adaptive ansatz grown from unitary coupled cluster theory to parametrize the ground states of the deuteron and 3He, compute their corresponding energies, and analyze the scaling of the required computational resources.
@@ I Introduction
In recent years, quantum computing has attracted considerable interest in nuclear theory. Given the limited coherence times of current quantum devices, research has focused on solving simple models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], developing computational methods and algorithms [18, 19, 20, 21, 22, 23, 24], tackling classically “hard” problems like nuclear scattering and dynamics [25, 26, 27, 28, 29, 30, 31, 32, 33], and gaining deeper insights into entanglement [34, 35, 36]. For recent reviews, we refer the reader to Refs. [37, 38, 39].
While this body of work reflects the excitement surrounding a new technology, more sober assessments [40, 41] raise doubts about whether a practical quantum advantage can be realized in computing ground-state energies. The skepticism stems from three main points. First, it is in general difficult to prepare initial states with a large overlap with the ground state. Second, many classical algorithms that scale polynomially with system size and are sufficient for accurate approximations of ground states. Finally, classical nuclear structure computations have advanced tremendously in scale and sophistication—beyond what was deemed possible just a decade ago [42, 43, 44, 45, 46].
Nevertheless, quantum algorithms offer a stringent theoretical advantage: they can provide guaranteed solutions within a specified error tolerance, something no classical method can ensure. This is typically achieved via Quantum Phase Estimation (QPE) [47] and its early fault-tolerant variants [48, 49, 50, 51]. The cost of QPE is determined by two main components: the preparation of an approximate initial state [52, 53] and the implementation of real-time evolution under the Hamiltonian [54, 55]. While fault-tolerant quantum computers remain under development, it is interesting to identify models well suited to quantum computation and to estimate the resources needed to solve them. In this paper, we address how to construct scalable nuclear Hamiltonians and prepare initial states using ADAPT-VQE [1, 56].
@@
However, in a Gray-code encoding (where a matrix representation of the $A$-body Hamiltonian is used ) [5, 16, 13] the number of Hamiltonian matrix elements increases faster than the Hilbert space dimension [57]. Thus, the translation of Hamiltonians into Gray-code operators scales exponentially with $A$. We see that in this framework quantum computing offers an exponential advantage for state representation while the construction and processing of the Hamiltonian operators remains exponentially costly. This is seen, e.g., in the quantum computing of 6Li of Ref. [13]. There, the Hilbert space with angular momentum projection $J_{z}=0$ required for the computation of $J^{\pi}=0^{+}$ states had dimension $D=10$ (see Table XIII of that work) and required 4 qubits, but the number of Gray-code operators was 134 (see Table XX of that work), which exceeded the number $D(D+1)/2=55$ of independent matrix elements of a $D\times D$ real-symmetric Hamiltonian matrix.
Instead, we work in the framework of second quantization. Then, $n_{q}$ qubits can represent the same number of single-particle states. Nuclear Hamiltonians consist of the kinetic energy (a one-body operator) and two- and three-body interactions. They are of the form
@@
We see that the Hamiltonian consists of ${\cal O}(n_{q}^{6})$ terms. Working in an axially symmetric framework (“$m$-scheme”) of the nuclear shell model [58] or in a momentum-space basis – where the total momentum is conserved [59] – reduces this number to about ${\cal O}(n_{q}^{5})$ and ${\cal O}(n_{q}^{3})$, respectively. While this is manageable in classical computing, such a scaling poses a significant effort on present-day quantum devices and might preclude quantum advantages [41]. As an example we again consider the quantum computation of 6Li, this time as performed in Ref. [7]. That approach used $n_{q}=12$ single-particle states (and qubits) but the Hamiltonian consisted of about 1000 Pauli terms. Hamiltonians in larger valence spaces that would be intractable on classical computers, such as the $pfsdg$-shell, easily have several hundred thousand Pauli terms. Clearly, one has to exploit the short range of the nuclear interaction. Using a lattice in position space [60, 61, 62, 63], the number of Hamiltonian terms is of order ${\cal O}(n_{q})$; it only grows linearly with the number of lattice sites because the nuclear interaction is short ranged.
@@
This paper is organized as follows: Section II introduces the lattice model space and Hamiltonian. In Sec. III we give details on the quantum computing algorithms used in Sec. IV to compute the structure of the deuteron and 3He. We conclude with a summary and perspectives in Sec. V. Some technical details are presented in the Appendices.
@@ II Lattice Hamiltonian
sites and use cubic boundary conditions. We label lattice sites using the indices $10,1,2,\ldots,n-$ 1. In general we need to employ $n_{q}=4n$ single-particle states because of two isospin and two spin projections. For the quantum computations we use $n_{q}$ qubits. We use a Hamiltonian from pionless effective field theory [66, 67]. At leading order, the interaction consists of a two-body contact and a three-body contact [68]. The Hamiltonian is
@@
where $m$ is the nucleon mass, $a$ is the lattice spacing, and $\mathbf{e}_{i}$ is a unit vector in the direction $i=x,y,z$. This is the leading order approximation of the Laplacian on the lattice. We use
@@
Any refinement, e.g., a more accurate kinetic energy or smeared contacts [69], would increase the number of Hamiltonian terms without changing this scaling as long as only short-range interactions are considered. It is clear that no other single-particle basis offers a more favorable scaling regarding the number of Hamiltonian terms. We adjust the coupling constants such that exact diagonalizations on sufficiently large lattices semi-quantatively reproduce the binding energies of light nuclei (see Appendix A for details). For the quantum computations in this work, we use $L=2$, $a=2.0\>\text{fm}$, $v=-9.0$, and $w=10.8$. On such a small lattice, finite-size corrections are substantial and the two- and three-body binding energies are about 12.9 and 29.5 MeV, respectively.
@@ III Quantum computation
The variational quantum eigensolver (VQE) has emerged as a promising approach for near-term quantum simulations, particularly for ground state problems in quantum many-body physics [1, 70]. The VQE approach combines a parametrized quantum circuit, known as an ansatz, with a classical optimizer to minimize the expectation value of the Hamiltonian. While conceptually appealing and implementable on current noisy hardware, standard VQE approaches face several practical limitations. These include a high number of measurements required to estimate expectation values with sufficient precision and the possibility of encountering barren plateaus-regions of vanishing gradients-in the optimization landscape [71].
To address some of these challenges, we employ the Adaptive Derivative-Assembled Pseudo-Trotter VQE (ADAPT-VQE) algorithm [56, 72], an iterative variant of VQE that dynamically constructs the variational ansatz during the optimization process. ADAPT-VQE adaptively grows the circuit by selecting the most impactful operators from a predefined operator pool. At each iteration, the operator that gives the steepest descent in energy is appended to the ansatz, ensuring that only the most relevant excitations are included. The advantage of ADAPT-VQE is that it is expected to deliver the most shallow circuit that can be expressed from the pool and at the same time accelerates the convergence of the variational optimization.
A crucial component of the ADAPT-VQE framework is the design of the operator pool, which must balance physical expressiveness and circuit efficiency. For our purposes, we construct the operator pool using the kinetic terms of the Hamiltonian (II) along with correlated two-body hopping operators. These operators are selected to preserve spin, isospin, and particle number. This physically informed construction helps reduce the search space and accelerates convergence while ensuring that the resulting variational ansatz remains compact and interpretable.
@@ III.1 System details
In this work, we perform quantum computations of the deuteron and 3He. The deuteron is the spin $S=1$ bound state of a proton and a neutron. This simplifies the calculation as follows. We choose an initial state
@@ III.2 ADAPT-VQE calculations
The intrinsic ground state must be invariant under discrete lattice translations. While it is in principle possible to generate initial states with this property (see Appendix C for details), we refrain from doing so for the following reasons. First, the unitary operator that yields a translationally invariant state when acting onto $|\phi_{0}\rangle$ will be approximated, e.g., via Trotterization. Thus, it cannot be implemented exactly. Second, the operators in the pool used by ADAPT-VQE break translational invariance. While one could contemplate using only translationally invariant combinations of operators, this would complicate the variation of the state. Third, the energy carried by spurious states decreases like $A^{-1}$ with increasing mass number and thus becomes less and less important in heavier nuclei. Finally we mention that one could also consider intrinsic Hamiltonians where the kinetic energy of the center of mass $T_{\rm CoM}$ is subtracted from Eq. (II). Then the number of Pauli terms in the Hamiltonian would increase from ${\cal O}(n_{q})$ to ${\cal O}(n_{q}^{2})$, because the $T_{\rm CoM}$ is a two-body operator in single-particle momenta.
@@ III.3 Sample complexity
where $a\in\mathbb{R}$ is a constant, $\hat{T}$ contains non-commuting terms, and $\hat{V}$ consists of mutually commuting terms that are simultaneously diagonalizable.
Let $\mathcal{T}$ and $\mathcal{V}$ denote the index sets corresponding to the non-commuting and commuting terms in $\hat{T}$ and $\hat{V}$, respectively. In the following, we will make the reasonable assumption that individual estimators from commuting observables remain uncorrelated when measured together. This allows us to group all the terms in $\mathcal{V}$ since they can measured simultaneously, resulting in a drastic shot reduction. The variance of the total energy estimate can be expressed as
@@
where $s_{V}$ is the shared number of measurements used to estimate all commuting terms, and $s_{\sigma}$ is the number of measurements allocated individually to each non-commuting term.
@@
$s_{V}$ $=\frac{1}{\varepsilon^{2}}\cdot\frac{\sum_{\sigma\in\mathcal{V}}v_{\sigma}^{2}}{\|\hat{H}\|_{2}^{2}}\,,$ (14)
$s_{\sigma}$ $=\frac{1}{\varepsilon^{2}}\cdot\frac{v_{\sigma}^{2}}{\|\hat{H}\|_{2}^{2}},\quad\text{for }\sigma\in\mathcal{T}\,,$ (15)
@@
This strategy ensures the most efficient distribution of samples for estimating $\langle\hat{H}\rangle$ up to error $\varepsilon$ by allocating a shared shot count to the commuting group $\mathcal{V}$ and individual shot counts to the non-commuting terms in $\mathcal{T}$.
Figure 1: Scaling of the number of samples $N_{\rm shots}$ required to reach a given precision in the lattice extent $L$ and the system size $A$. Going from $\varepsilon=1~{}\mathrm{MeV}$ (filled-in circles) to 0.1 MeV (open circles) requires two orders of magnitude more samples. For $A=2,3$ we are able to make simplifications such that ADAPT-VQE computations only require $n_{q}=2L^{3},3L^{3}$ qubits, respectively, rather than the general $n_{q}=4L^{3}$.
Based on our selected dimensionless couplings $v=-9.0$, $w=10.8$ and our lattice spacing $a=2.0\>\text{fm}$, we compute the shot budget required to estimate the energy within 1 MeV and $0.1$ MeV as a function of the lattice extent $L$ and system size $A$. The results, shown in Fig. 1, show that the number of shots required strongly depends on the desired precision. Going from $\varepsilon=1\>\mathrm{MeV}$ to $0.1\>\mathrm{MeV}$ requires approximately 100 times as many samples. For increasing lattice sizes the required number of samples scales like the volume $L^{3}$. For lattice sizes $L=2$–4, the number of shots required to reach $\varepsilon=1\>\mathrm{MeV}$ is on the order of $10^{4}$$10^{6}$ depending on the system, while increasing to $L=10$ increases this by another order of magnitude. Quantum computations with $N_{\rm shots}\sim 10^{6}$ are expensive, but possible on current hardware, and this is lower than the number of shots required for quantum chemistry applications, where as many as $10^{10}$ shots may be required to reach chemical accuracy in the case of, e.g., ethanol [78]. We note that those resource estimates are only for the resources required to estimate the energy of a given state and thus do not entail the optimization cost.

### v2

# Toward scalable quantum computations of atomic nuclei

@@ Abstract
We solve the nuclear two-body and three-body bound states via quantum simulations of pionless effective field theory on a lattice in position space. While the employed lattice remains small, the usage of local Hamiltonians including two- and three-body forces ensures that the number of Pauli terms scales linearly with increasing numbers of lattice sites. We use an adaptive ansatz grown from unitary coupled cluster theory to parametrize the ground states of the deuteron and 3He, compute their corresponding energies, and analyze the scaling of the required computational resources. Our quantum simulations reproduce exact benchmarks for 2H and 3He within 100 keV, requiring at most 30 layers in the ansatz and thus resulting in modest circuit depths. Additionally, we find the number of shots required to reach a given precision scales linearly in the lattice size and more mildly in the system size. Based on the agreement with exact benchmarks and mild scaling, we conclude that this can be an efficient, scalable approach for quantum computations of nuclear ground states, particularly to prepare initial states for quantum phase estimation or other filtering algorithms.
@@ I Introduction
In recent years, quantum computing has attracted considerable interest in nuclear theory. Given the limited coherence times of current quantum devices, research has focused on solving nuclear models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], developing computational methods and algorithms [20, 21, 22, 23, 24, 25, 26], tackling classically “hard” problems like nuclear scattering and dynamics [27, 28, 29, 30, 31, 32, 33, 34, 35, 36], and gaining deeper insights into entanglement [37, 38, 39]. For recent reviews, we refer the reader to Refs. [40, 41, 42].
While this body of work reflects the excitement surrounding a new technology, more sober assessments [43, 44] raise doubts about whether a practical quantum advantage can be realized in computing ground-state energies. The skepticism stems from two main points. First, it is in general difficult to prepare initial states with a large overlap with the ground state. Second, many classical algorithms scale polynomially with system size and are sufficiently accurate to compute various properties of nuclear ground states [45, 46, 47, 48, 49, 50]. Such computations have advanced tremendously in scale and sophistication—beyond what was deemed possible just a decade ago.
Nevertheless, quantum algorithms offer a stringent theoretical advantage: They can provide guaranteed solutions within a specified error tolerance, something no classical method can ensure. This is typically achieved via quantum phase estimation (QPE) [51] and its early fault-tolerant variants [52, 53, 54, 55]. The cost of QPE is determined by two main components: the preparation of an approximate initial state [56, 57] and the implementation of real-time evolution under the Hamiltonian [58, 59]. While fault-tolerant quantum computers remain under development, it is interesting to identify models well suited to quantum computation and to estimate the resources needed to solve them. In this paper, we address how to construct scalable nuclear Hamiltonians and prepare initial states using the adaptive derivative-assembled pseudo-Trotter variational quantum eigensolver (ADAPT-VQE) [1, 60].
@@
However, in a Gray-code encoding (where a matrix representation of the $A$-body Hamiltonian is used) [5, 16, 13, 61, 62, 36, 19] the number of Hamiltonian matrix elements increases faster than the Hilbert space dimension [63]. Thus, the translation of Hamiltonians into Gray-code operators scales exponentially with $A$. We see that in this framework quantum computing offers an exponential advantage for state representation while the construction and processing of the Hamiltonian operators remains exponentially costly. This is seen, e.g., in the quantum computing of 6Li of Ref. [13]. There, the Hilbert space with angular momentum projection $J_{z}=0$ required for the computation of $J^{\pi}=0^{+}$ states had dimension $D=10$ (see Table XIII of that work for a complete list of $J^{\pi}=0^{+}$ states) and required four qubits, but the number of Gray-code operators was 134 (see Table XX of that work for a complete list of Gray-code operators), which exceeded the number $D(D+1)/2=55$ of independent matrix elements of a $D\times D$ real-symmetric Hamiltonian matrix.
Instead, we work in the framework of second quantization, as is also used in, e.g., Refs. [6, 10, 17, 15]. Then, $n_{q}$ qubits can represent the same number of single-particle states. Nuclear Hamiltonians consist of the kinetic energy (a one-body operator) and two- and three-body interactions. They are of the form
@@
We see that the Hamiltonian consists of ${O}(n_{q}^{6})$ terms. Working in an axially symmetric framework (“$m$-scheme”) of the nuclear shell model [64] or in a momentum-space basis – where the total momentum is conserved [65] – reduces this number to about ${O}(n_{q}^{5})$ and ${O}(n_{q}^{3})$, respectively. While this is manageable in classical computing, such a scaling poses a significant effort on present-day quantum devices and might preclude quantum advantages [44]. As an example we again consider the quantum computation of 6Li, this time as performed in Ref. [7]. That approach used $n_{q}=12$ single-particle states (and qubits) but the Hamiltonian consisted of about 1000 Pauli terms. Hamiltonians in larger valence spaces that would be intractable on classical computers, such as the pfsdgpfsdg shell, easily have several hundred thousand Pauli terms. Clearly, one has to exploit the short range of the nuclear interaction. Using a lattice in position space [66, 67, 68, 69], the number of Hamiltonian terms is of order ${O}(n_{q})$; it only grows linearly with the number of lattice sites because the nuclear interaction is short ranged.
@@
This paper is organized as follows: Section II introduces the lattice model space and Hamiltonian. In Sec. III we give details on the quantum computing algorithms used in Sec. IV to compute the structure of the deuteron and 3He. We conclude with a summary and perspectives in Sec. V. Some technical details are presented in the appendixes.
@@ II Lattice Hamiltonian
sites and use cubic boundary conditions. We label lattice sites using the indices $0,1,2,\ldots,n-1$. In general we need to employ $n_{q}=4n$ single-particle states because of two isospin and two spin projections. For the quantum computations we use $n_{q}$ qubits. We use a Hamiltonian from pionless effective field theory (EFT) [71, 72]. At leading order, the interaction consists of a two-body contact and a three-body contact [73]. The Hamiltonian is
@@
where $m$ is the nucleon mass, $a$ is the lattice spacing, and $\mathbf{e}_{i}$ is a unit vector in the direction $i=x,y,z$. This is the leading order approximation of the Laplacian on the lattice and corrections are of order ${O}(a^{2})$ relative to the leading approximation (4). We use
@@
The Hamiltonian (II) implements pionless EFT at leading order. That theory is a low-energy effective field theory of quantum chromodynamics [71] and has been used to describe few-nucleon systems [73, 74, 75, 76, 77]. It reflects that nuclear interactions – in their crudest approximation – are short ranged and nearly spin-isospin symmetric.
@@
The Hamiltonian (II) may be systematically improved for more precise computations of nuclei using techniques from effective field theory [72]. More accurate lattice Hamiltonians can be found in Refs. [66, 78, 48, 69] and we briefly mention some of their features. Improved approximations of the Laplacian couple lattice sites up to next-to-next-to-nearest neighbors [79]. A more accurate description of nuclei requires higher-order contributions, nonlocal interactions [78] and interactions from pion-exchange [80], which both introduce spin-orbit and tensor interactions. Higher-order contact interactions have a short, but finite range. Pion-exchange has the range of the inverse pion mass $m_{\pi}^{-1}\approx 1.5$ fm and will predominantly couple nearest and next-to-nearest neighbors for typically used lattice spacings $a\sim 1.5$22 fm.
Crucially, any such refinement involving finite-range interactions does not change the fact that the number of matrix elements scales linearly in the basis size $n_{q}$. This also applies to finite-range approximations of the Laplacian. [^1] The exact kinetic energy is local in Fourier space and thus completely nonlocal in position space, which would give a scaling of $O(n_{q}^{2})$. In the case that finite-range approximations to the Laplacian are insufficient for the desired precision, this scaling would still be favorable to the $O(n_{q}^{3})$$O(n_{q}^{5})$ scaling in other bases. This can be easily understood. The finite range allows particles to interact in a small volume, but this volume is independent of the total volume of the simulation. The result is that the scaling is still linear in the basis size, simply with a different prefactor. This prefactor may be large, and this will accordingly increase the computational cost of the simulation, but for computations in large bases it is still very favorable compared with $O(n_{q}^{3})$ and $O(n_{q}^{5})$ scaling in other bases. (We recall that state of the art ab initio calculations of nuclei on classical computers employ basis sizes using 2000–5000 single-particle states [47, 48].) It is clear that no other single-particle basis offers a more favorable scaling regarding the number of Hamiltonian terms.
The Coulomb interaction is long range and delivers small, but important contributions for nuclear structure. In pionless EFT, it contributes at next-to-leading order for ground-state energies [81, 82, 83] and may be treated in perturbation theory [79] following a nonperturbative ground-state solution. Such a perturbative treatment would also be an option in quantum computing contexts, with the obvious challenge that the number of nonzero matrix elements scales like $O(n_{q}^{2})$ (due to being long range and local). Efficiently evaluating this contribution either via classical postprocessing or via appropriate quantum algorithms will be an important problem for future work.
We adjust the coupling constants such that exact diagonalizations on sufficiently large lattices semiquantitatively reproduce the binding energies of light nuclei (see Appendix A for details). For the quantum computations in this work, we use $L=2$, $a=2.0\>\text{fm}$, v=9.0v=-9.0, and $w=10.8$. On such a small lattice, finite-size corrections are substantial and the two- and three-body binding energies are about 12.9 and 29.5 MeV, respectively.
@@
Let us briefly comment on the special structure of the Hamiltonian (II) where the potential is diagonal in the chosen lattice basis and the kinetic energy is off-diagonal. The recent study [84] revealed that Hartree Fock yields about 81% of the ground-state energy for 3He computed with a lattice spacing $a=2$ fm (as used in this work). Furthermore, the coupled-cluster method with single and double excitations [85, 86] or the in-medium similarity renormalization group [87, 88] yield little improvement over the Hartree-Fock energy. Our results below meet benchmarks from exact diagonalization, and this shows that three-particle–three-hole excitations are fully captured (for the 3He nucleus) in our quantum computations. In summary, Hartree Fock captures a large portion of the ground-state energy because the kinetic energy is nonlocal while two- and three-body interactions are on site. However, obtaining the exact energy requires exact or high-order many-body expansion methods.
@@ III Quantum computation
The variational quantum eigensolver (VQE) has emerged as a promising approach for near-term quantum simulations, particularly for ground state problems in quantum many-body physics [1, 89]. The VQE approach combines a parametrized quantum circuit, known as an ansatz, with a classical optimizer to minimize the expectation value of the Hamiltonian. While conceptually appealing and implementable on current noisy hardware, standard VQE approaches face severe practical limitations. These include a high number of measurements required to estimate expectation values with sufficient precision and the possibility of encountering barren plateaus, i.e., regions of vanishing gradients, in the optimization landscape [90, 91].
To mitigate the trainability issue, we employ the adaptive derivative-assembled pseudo-Trotter VQE (ADAPT-VQE) algorithm [60, 92], an iterative variant of VQE that dynamically constructs the variational ansatz during the optimization process. ADAPT-VQE adaptively grows the circuit by selecting the most impactful operators from a predefined operator pool. At each iteration, the operator that gives the steepest descent in energy is appended to the ansatz, ensuring that only the most relevant excitations are included. The advantage of ADAPT-VQE is that it is expected to deliver the most shallow circuit that can be expressed from the pool and at the same time accelerates the convergence of the variational optimization. ADAPT-VQE has been applied in nuclear physics, albeit mainly within the shell-model framework [7, 10]. In this work, we focus instead on a scalable lattice formulation, which provides a complementary setting for exploring ADAPT-VQE.
A crucial component of the ADAPT-VQE framework is the design of the operator pool, which must balance physical expressiveness and circuit efficiency. For our purposes, we construct the operator pool using the kinetic terms of the Hamiltonian (II) along with correlated two-body hopping operators. These operators are selected to preserve spin, isospin, and particle number. This physically informed construction helps reduce the search space and accelerates convergence while ensuring that the resulting variational ansatz remains compact and interpretable. One can imagine further tailoring the operator pool the problem of interest to efficiently capture the expected dominant states. Such tailoring, if done correctly, can lead to shallower circuits in VQE computations, as was shown in Ref. [15] for shell-model quantum computations of 58Ni. Developing an expressive, but general operator pool is clearly important for generally parametrizing the wave function while minimizing the circuit depth obtained using variational algorithms.
@@ III.1 System details
In this work, we perform quantum computations of the deuteron and 3He. The deuteron is the spin $S=1$ bound state of a proton and a neutron. This simplifies the calculation as follows: We choose an initial state
@@ III.2 ADAPT-VQE calculations
The intrinsic ground state must be invariant under discrete lattice translations. While it is in principle possible to generate initial states with this property (see Appendix C for details), we refrain from doing so for the following reasons. First, the unitary operator that yields a translationally invariant state when acting on $|\phi_{0}\rangle$ will be approximated, e.g., via Trotterization. Thus, it cannot be implemented exactly. Second, the operators in the pool used by ADAPT-VQE break translational invariance. While one could contemplate using only translationally invariant combinations of operators, this would complicate the variation of the state. Third, the energy carried by spurious states decreases like $A^{-1}$ with increasing mass number and thus becomes less and less important in heavier nuclei. Finally we mention that one could also consider intrinsic Hamiltonians where the kinetic energy of the center of mass $T_{\rm CoM}$ is subtracted from Eq. (II). Then the number of Pauli terms in the Hamiltonian would increase from ${O}(n_{q})$ to ${O}(n_{q}^{2})$, because the $T_{\rm CoM}$ is a two-body operator in single-particle momenta.
@@ III.3 Sample complexity
where $a\in\mathbb{R}$ is a constant, $\hat{T}$ contains noncommuting terms, and $\hat{V}$ consists of mutually commuting terms that are simultaneously diagonalizable.
Let $\mathcal{T}$ and $\mathcal{V}$ denote the index sets corresponding to the noncommuting and commuting terms in $\hat{T}$ and $\hat{V}$, respectively. In the following, we make the reasonable assumption that individual estimators from commuting observables remain uncorrelated when measured together. This allows us to group all the terms in $\mathcal{V}$ since they can measured simultaneously, resulting in a drastic shot reduction. The variance of the total-energy estimate can be expressed as
@@
where $s_{V}$ is the shared number of measurements used to estimate all commuting terms, and $s_{\sigma}$ is the number of measurements allocated individually to each noncommuting term.
@@
$s_{V}$ $=\frac{1}{\varepsilon^{2}}\frac{\sum_{\sigma\in\mathcal{V}}v_{\sigma}^{2}}{\|\hat{H}\|_{2}^{2}}\,$, (14)
$s_{\sigma}$ $=\frac{1}{\varepsilon^{2}}\frac{v_{\sigma}^{2}}{\|\hat{H}\|_{2}^{2}}\quad\text{for }\sigma\in\mathcal{T}\,$, (15)
@@
This strategy ensures the most efficient distribution of samples for estimating $\langle\hat{H}\rangle$ up to error $\varepsilon$ by allocating a shared shot count to the commuting group $\mathcal{V}$ and individual shot counts to the noncommuting terms in $\mathcal{T}$.
Figure 1: Scaling of the number of samples $N_{\rm shots}$ required to reach a given precision in the lattice extent $L$ and the system size $A$. Going from $\varepsilon=1~\mathrm{MeV}$ (filled circles) to 0.1 MeV (open circles) requires two orders of magnitude more samples. For $A=2,3$ we are able to make simplifications such that ADAPT-VQE computations only require $n_{q}=2L^{3},3L^{3}$ qubits, respectively, rather than the general $n_{q}=4L^{3}$.
Based on our selected dimensionless couplings v=9.0v=-9.0, $w=10.8$ and our lattice spacing $a=2.0\>\text{fm}$, we compute the shot budget required to estimate the energy within 1 and $0.1$ MeV as a function of the lattice extent $L$ and system size $A$. The results, shown in Fig. 1, show that the number of shots required strongly depends on the desired precision. Going from $\varepsilon=1$ to $0.1\>\mathrm{MeV}$ requires approximately 100 times as many samples. For increasing lattice sizes the required number of samples scales like the volume $L^{3}$. For lattice sizes L=2L=2–4, the number of shots required to reach $\varepsilon=1\>\mathrm{MeV}$ is on the order of $10^{4}$$10^{6}$ depending on the system, while increasing to $L=10$ increases this by another order of magnitude. Quantum computations with $N_{\rm shots}\sim 10^{6}$ are expensive, but possible on current hardware, and this is lower than the number of shots required for quantum chemistry applications, where as many as $10^{10}$ shots may be required to reach chemical accuracy in the case of, e.g., ethanol [98]. We note that those resource estimates are only for the resources required to estimate the energy of a given state and thus do not entail the optimization cost.